# Gaussian Variational Inference with Non-Gaussian Factors for State Estimation: A UWB Localization Case Study

### v1

@@ I Introduction
Autonomous robots must estimate their position, orientation, and other key pieces of information to operate effectively. This collection of dynamic variables, known as the state of the robot, fully describes the system at a given time. Consistent and accurate state estimation is essential for high-level decision-making and the support of downstream tasks such as guidance and control. Within state estimation, approaches to localization can be broadly split into the filtering and batch approaches [Barfoot_2024]. Filtering methods use data up to the current timestep to estimate the robot’s state, whereas batch methods use all available data over a given time period. The latter forms the foundation of this work. Taking a Bayesian perspective, the goal of state estimation is to compute the state $\mbf{x}\in\mathbb{R}^{n_{x}}$, given a set of measurements $\mbf{y}\in\mathbb{R}^{n_{y}}$, that is, the full posterior $p(\mbf{x}|\mbf{y})$ is computed, given a prior $p(\mbf{x})$. Using Bayes’ rule, the relationship
@@
is exploited to compute the posterior. For nonlinear measurement models $p(\mbf{y}|\mbf{x})$, the posterior is not a Gaussian probability density function (PDF). Currently, the majority of the batch state estimation literature is based on solving a maximum-a-posteriori (MAP) problem, which involves computing the maximum, that is the mode, of the Bayesian posterior. An alternative to MAP is variational inference [Bishop_2006], which approximates the posterior by fitting a distribution that minimizes the Kullback–Leibler (KL) divergence. In [esgvi_paper], exactly sparse Gaussian variational inference (ESGVI) was introduced as a tractable method for computing Gaussian approximations of the posterior in large-scale nonlinear batch state estimation problems, by exploiting the sparsity associated with robotics problems. By accounting for the full shape of the posterior distribution, ESGVI was shown to outperform MAP in nonlinear state estimation problems. However, in its current form, ESGVI operates on vector-valued states and, although the use of a robust cost function was suggested as a future line of inquiry in [esgvi_paper], it has only been evaluated with Gaussian-distributed sensor models. In practice, sensor measurements can deviate significantly from Gaussian assumptions, producing outliers that can severely degrade estimation performance. This issue is particularly evident in ultra-wideband (UWB) localization, where ranging measurements are increasingly used in robotics due to their long-range and low-cost characteristics [Shalaby_Cossette_Forbes_Le_Ny_2023]. Nevertheless, UWB measurements are often corrupted by non-line-of-sight (NLOS) effects and multipath propagation, leading to heavy-tailed and asymmetric noise distributions that must be accounted for [Kim_Yoon_Lee_2024]. To address these limitations, ESGVI is extended to operate on matrix Lie group (MLG) states, a modest but necessary step, as MLGs naturally capture the geometry of states commonly found in robot navigation problems [Solà_Deray_Atchuthan_2021]. In addition, non-Gaussian measurement factors are incorporated, providing a more flexible way to model heavy-tailed and skewed noise distributions. While such factors have been explored in related estimation frameworks [Vilà-Valls_Vincent_Closas_2019], the contribution here lies in demonstrating their integration within the ESGVI setting and evaluating their impact in practice. Although the proposed extensions are general, their effectiveness is illustrated through a specific UWB example for concreteness and ease of interpretation. Building on [esgvi_paper], the contributions of this letter are as follows:
@@
  • Evaluating ESGVI in a real-world, outlier rich, UWB localization experiment, demonstrating greater accuracy and consistency.
@@ III-A Probabilistic Model Specification
While attention is restricted to vector states here for clarity, in later sections this framework will be generalized to accommodate the MLG state, $\mbf{X}_{k}$. The state evolves over time according to a stochastic state transition with the Markov property, i.e., only dependent on the previous state [Sarkka, Property 4.1]. It is modelled as a Gaussian distribution,
@@ III-B State of the Art
By assuming Gaussian noise and minimizing the negative logarithm of the posterior, the MAP estimate can be formulated as a nonlinear least-squares (NLS) problem [Barfoot_2024, §3.1.2]. Expressed in this sense, the estimator minimizes a sum-of-squared-errors cost function, which makes it highly sensitive to outliers [Barfoot_2024, §5.4]. To address the sensitivity of MAP-based NLS estimators to outliers, a variety of strategies have been explored, particularly in UWB-based state estimation. For example, Omni-Swarm [omni_swarm] introduced a visual–inertial–UWB estimation pipeline formulated as a MAP-based NLS optimization problem, where UWB outliers were rejected using fixed elevation-angle criteria. Distributed approaches [Fishberg_How_2022, murp, vir_slam] extend this concept to multi-agent settings, also solving MAP NLS problems while rejecting inconsistent range measurements through averaging, geometric or max velocity constraints. More recently, learning-based methods have been employed to detect and discard NLOS UWB measurements. In [Qi_Zhao_Guo_Kaliuzhnyi_Wang_2025] and [Jiang_Shen_Chen_Chen_Liu_Bo_2020], convolutional neural network (CNN) and long short-term memory (LSTM) architectures were proposed to classify and reject suspected NLOS observations. However, since no model can guarantee perfect outlier identification, residual contamination typically persists in the dataset, introducing skewed measurement noise.
To more systematically account for these residual effects, M-estimators incorporate robust cost functions directly within the NLS formulation, allowing the optimization to down-weight the influence of large residuals rather than reject measurements outright. Various robust loss functions have been proposed [zhang1997parameter], and in the context of UWB localization, several noise parameterizations have been studied. For example, in [uvip] a Huber loss was integrated into the NLS optimization, while [Shalaby_Cossette_Forbes_Le_Ny_2023] proposed a Cauchy loss, however these symmetric losses ignore the asymmetry present in NLOS noise. To capture skewed characteristics, [Kok_Hol_Schön_2015, Kim_Yoon_Lee_2024] used an asymmetric two-piece distribution. A similar asymmetric Cauchy distribution is used in this letter defined by
@@
where $r(\cdot,\cdot)$ is the residual, $c^{-}$ and $c^{+}$ are scale parameters, and $\alpha=2/[\pi(c^{+}+c^{-})]$ normalizes the distribution. The asymmetric Cauchy loss provides outlier robustness while explicitly modeling skewed measurement noise. A secondary flexible alternative for modeling skewed distributions is to employ a Gaussian mixture measurement model [Zhao_Goudar_Tang_Qiao_Schoellig_2023],
@@
with weights $w_{j}\geq 0,\ \sum_{j}w_{j}=1$ and $\bm{\psi}_{j}=(\mu_{j},\Sigma_{j})$ [Pfeifer_Lange_Protzel_2021]. Component weights and parameters are calibrated separately, or solved for in parallel with the state $\mbf{x}$, in an Expectation Maximization (EM) framework [Bishop_2006, §9.3]. The Gaussian mixture measurement model of (10) is incorporated into the joint distribution of (7), and MAP estimation (8) produces a state estimate. Although Gaussian mixture models (GMMs) are flexible enough to capture skewed noise distributions, they require additional parameters and, while incorporation into a NLS framework is possible, it demands careful handling [Pfeifer_Lange_Protzel_2021]. Moreover, a MAP estimator provides only a point estimate of the state, with uncertainty obtained using an ad-hoc Laplace approximation [Barfoot_2024, §4.3.1]. Alternatively, while the methods discussed thus far rely on MAP-based optimization, ESGVI offers a fundamentally different probabilistic framework that estimates both the state and its uncertainty directly through variational inference. Performing parameter learning in parallel with ESGVI [Wong_Yoon_Schoellig_Barfoot_2020] has been shown to improve robustness, as jointly estimating the covariance with an appropriate prior implicitly rejects outliers. While robustness was identified as an open direction in [esgvi_paper], measurement models specific to a non-Gaussian distribution, such as the Skew-Laplace in (6) have not yet been studied directly within the ESGVI framework. Consequently, in this work, comparisons are made against robust MAP approaches rather than existing ESGVI methods. In the following, ESGVI is shown to naturally incorporate robust, non-Gaussian measurement models with matrix Lie group states, enabling principled estimation under heavy-tailed or asymmetric noise. This is illustrated using a Skew-Laplace model for UWB localization, capturing NLOS-induced asymmetry.
@@ IV Gaussian Variational Inference
As introduced previously, an alternative to MAP estimation is Gaussian variational inference (GVI), which fits a Gaussian approximation $q(\mbf{x})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, to the posterior $p(\mbf{x|y,u})$, by minimizing the Kullback-Leibler (KL) divergence [Bishop_2006],
@@
where $\phi(\mbf{x})=-\ln p(\mbf{x,y,u})$ is the negative log joint state and measurement likelihood, and $|\bm{\Sigma}^{-1}|$ is the determinant of the information matrix of the variational approximation, $q(\mbf{x})$. In the work of [esgvi_paper], an exactly sparse Gaussian variational inference (ESGVI) optimization scheme was defined for the vector state $\mbf{x}$, minimizing $V(q)$. The iterative update to the mean $\bm{\mu}$, and information matrix $\bm{\Sigma}^{-1}$, are defined according to
@@
which reduces the computation to smaller, factor-specific subsets. Next, using Stein’s lemma [Barfoot_2024, §2.2.16], it was demonstrated how taking derivatives of $\phi_{k}(\cdot)$ can be avoided [esgvi_paper],
@@
Lastly, these marginal expectations can be approximated using a cubature method [Sarkka, §6.3], that is
@@ IV-A ESGVI with Matrix Lie Group States
Up to this point, the discussion has been restricted to vector-valued states $\mbf{x}_{k}$, for simplicity. However, as introduced previously, some robotics states are more naturally expressed as MLG states $\mbf{X}_{k}$. As such the variational algorithm (13) must be reformulated to work with MLG states. To do so, uncertainty must be defined in a way that is compatible with the underlying group structure [Barfoot_Furgale_2014]. A Gaussian distribution associated with a MLG state is defined by perturbing the mean element $\bar{\mbf{X}}$ with a zero-mean Gaussian random variable $\delta\bm{\xi}$. Formally,
@@
Although the expressions appear structurally similar to their Euclidean counterparts, their interpretation on MLGs requires further clarification, provided in Appendix A-B. Numerically, the expectations of (18) are evaluated using a cubature method [Sarkka, §6.3] such that
@@
The flexibility of $\phi_{k}(\mbf{X}_{k})$ to represent any likelihood model, including Huber or Cauchy, was noted in [esgvi_paper], and such factors are incorporated into ESGVI using the same procedure described above. Bringing these components together, the generalized form of (13) for MLGs is reformulated as,
@@ V Simulation and Experimental Results
The proposed method is validated in simulation and in experiments, and compared against two robust on-manifold MAP baselines, one employing the asymmetric Cauchy loss (MAP-C) shown in (9) and the other a GMM (MAP-GMM) to model skewed measurements. All estimators are implemented in Python within a common factor-graph-based framework. No comparison is made against a vector-valued ESGVI formulation, as the advantages of MLG-based state representations, such as improved consistency and avoidance of singularities, are well established in the literature [long2013banana, Barfoot_Furgale_2014]. The following sections present results first on simulated data, which allows controlled evaluation of performance, and then on real-world experiments to demonstrate practical applicability.
@@ V-A Simulation Results
where $\hat{\mbf{X}}_{i,k}$, and $\mbf{X}_{i,k}$ are the estimated and ground-truth states, for trajectory $i$, at time $t_{k}$, with RMSE reported separately for rotation ($\mathrm{r}\mathrm{a}\mathrm{d}$) and translation ($\mathrm{m}$). Consistency, meaning the ability of the estimator to correctly quantify its own uncertainty, is measured using the average normalized estimation error squared (aNEES) [bar2001estimation, §5.4],
@@
where a perfectly consistent estimator has an aNEES equal to one, due to the normalization in (28) by the dimension of the state $n_{x}$. Performance comparisons are based on $50$ Monte Carlo trials over a trajectory of $400$ poses, with measurements and initial conditions regenerated randomly in each trial, with odometry being produced at $100$ ($\mathrm{H}\mathrm{z}$), and range measurements occurring at $10$ ($\mathrm{H}\mathrm{z}$). In all trials, the same GMM and Skew-Laplace parameters are used, fit using 5,0005{,}000 samples as shown in Figure 1(b). All algorithms are initialized with dead-reckoned gyroscope and wheel odometry measurements, though ESGVI can optionally be “warm-started” using a MAP-based solution to accelerate convergence. Finally, throughout this section, ESGVI uses a third-order Gauss-Hermite cubature method [Sarkka, §6.3] to evaluate the expectations in (19). As seen in Table I, all three estimators have similar performance. The GMM model is the most consistent, while the RMSE is comparable between the ESGVI and the MAP estimators. ESGVI is slightly overconfident compared to the MAP estimators, which follows the results of [Bishop_2006] as minimizing the KL-divergence, $\textrm{KL}(q||p)$ can result in a Gaussian that is too confident, underestimating the variance.
@@ V-B Experimental Setup
Experiments were conducted using the Clearpath Husky vehicle, equipped with four DWM1000 Decawave UWB tags, as shown in Figure 2. These tags measure ranges to a set of fixed UWB anchors positioned throughout the environment, which are then used for localization. UWB ranging relies on measuring the time-of-flight of radio signals [Shalaby_Cossette_Forbes_Le_Ny_2023] and is therefore sensitive to non-line-of-sight (NLOS) propagation and multipath effects, which can introduce bias in the measured ranges, as illustrated in Figure 1. Such conditions are common in indoor environments, where obstacles, clutter, and frequent foot traffic create dynamic obstructions that lead to skewed measurements. To reproduce these effects in a controlled manner, obstacles made of metal, wood, and foam were placed near selected UWB anchors, as shown in Figure 2.
@@
The Clearpath Husky was driven along ten independent trajectories designed to randomly expose the UWB range measurements to alternating LOS and NLOS conditions. An example of the sixth trajectory is shown in Figure 3. Ground-truth poses were recorded at $120$ ($\mathrm{H}\mathrm{z}$) using a Vicon motion capture system and used for estimator performance evaluation. The robot was operated at a nominal forward velocity of $0.4$ ($\mathrm{m}\mathrm{/}\mathrm{s}$), resulting in trajectories of approximately $60$ seconds each in duration within a $40$ ($\mathrm{m}^{2}$) indoor test area. UWB range measurements were collected at $30$ ($\mathrm{H}\mathrm{z}$) per tag, yielding over 75,00075{,}000 range observations across all ten trajectories. In addition to NLOS effects, UWB measurements are affected by other error sources, including antenna delays, irregular radiation patterns, and clock skews between UWB tags. To account for these factors, the calibration procedure described in [Shalaby_Cossette_Forbes_Le_Ny_2023] was applied. In particular, the antenna-delay calibration was performed, and the data-driven model relating received signal power to attitude-dependent range bias was reused to isolate for NLOS-induced bias effects. Range measurement errors were then computed with respect to the ground-truth poses, across all trajectories. The resulting error distribution, shown in Figure 4, clearly demonstrates the influence of obstacles through its pronounced skewed tails. Based on this distribution, the asymmetric Cauchy, GMM, and Skew-Laplace models were fitted. The GMM utilized three mixture components, consistent with the methodology of [Zhao_Goudar_Tang_Qiao_Schoellig_2023], where UWB errors were modeled in a similar fashion. All three models effectively capture the skewness arising from the prevalence of NLOS measurements, with the Skew-Laplace model exhibiting the most conservative behavior for highly biased ranges, while the asymmetric Cauchy and GMM models more closely align with the empirical histogram. Notably, the experimental error distribution closely mirrors the simulated case shown in Figure 1(b), reinforcing the validity of the simulation assumptions.

### v2

# IEEE Xplore: https://ieeexplore.ieee.org/document/11347486

@@ I Introduction
Autonomous robots must estimate their position, orientation, and other key pieces of information to operate effectively. This collection of dynamic variables, known as the state of the robot, fully describes the system at a given time. Consistent and accurate state estimation is essential for high-level decision-making and the support of downstream tasks such as guidance and control. Within state estimation, approaches to localization can be broadly split into the filtering and batch approaches [4]. Filtering methods use data up to the current timestep to estimate the robot’s state, whereas batch methods use all available data over a given time period. The latter forms the foundation of this work. Taking a Bayesian perspective, the goal of state estimation is to compute the state $\mbf{x}\in\mathbb{R}^{n_{x}}$, given a set of measurements $\mbf{y}\in\mathbb{R}^{n_{y}}$, that is, the full posterior $p(\mbf{x}|\mbf{y})$ is computed, given a prior $p(\mbf{x})$. Using Bayes’ rule, the relationship
@@
is exploited to compute the posterior. For nonlinear measurement models $p(\mbf{y}|\mbf{x})$, the posterior is not a Gaussian probability density function (PDF). Currently, the majority of the batch state estimation literature is based on solving a maximum-a-posteriori (MAP) problem, which involves computing the maximum, that is the mode, of the Bayesian posterior. An alternative to MAP is variational inference [5], which approximates the posterior by fitting a distribution that minimizes the Kullback–Leibler (KL) divergence. In [2], exactly sparse Gaussian variational inference (ESGVI) was introduced as a tractable method for computing Gaussian approximations of the posterior in large-scale nonlinear batch state estimation problems, by exploiting the sparsity associated with robotics problems. By accounting for the full shape of the posterior distribution, ESGVI was shown to outperform MAP in nonlinear state estimation problems. However, in its current form, ESGVI operates on vector-valued states and, although the use of a robust cost function was suggested as a future line of inquiry in [2], it has only been evaluated with Gaussian-distributed sensor models. In practice, sensor measurements can deviate significantly from Gaussian assumptions, producing outliers that can severely degrade estimation performance. This issue is particularly evident in ultra-wideband (UWB) localization, where ranging measurements are increasingly used in robotics due to their long-range and low-cost characteristics [15]. Nevertheless, UWB measurements are often corrupted by non-line-of-sight (NLOS) effects and multipath propagation, leading to heavy-tailed and asymmetric noise distributions that must be accounted for [10]. To address these limitations, ESGVI is extended to operate on matrix Lie group (MLG) states, a modest but necessary step, as MLGs naturally capture the geometry of states commonly found in robot navigation problems [Solà_Deray_Atchuthan_2021]. In addition, non-Gaussian measurement factors are incorporated, providing a more flexible way to model heavy-tailed and skewed noise distributions. While such factors have been explored in related estimation frameworks [Vilà-Valls_Vincent_Closas_2019], the contribution here lies in demonstrating their integration within the ESGVI setting and evaluating their impact in practice. Although the proposed extensions are general, their effectiveness is illustrated through a specific UWB example for concreteness and ease of interpretation. Building on [2], the contributions of this letter are as follows:
@@
  • Evaluating ESGVI in a real-world, outlier rich, UWB localization experiment, demonstrating greater accuracy and comparable consistency.
@@ III-A Probabilistic Model Specification
While attention is restricted to vector states here for clarity, in later sections this framework will be generalized to accommodate the MLG state, $\mbf{X}_{k}$. The state evolves over time according to a stochastic state transition with the Markov property, i.e., only dependent on the previous state [14, Property 4.1]. It is modelled as a Gaussian distribution,
@@ III-B State of the Art
By assuming Gaussian noise and minimizing the negative logarithm of the posterior, the MAP estimate can be formulated as a nonlinear least-squares (NLS) problem [4, §3.1.2]. Expressed in this sense, the estimator minimizes a sum-of-squared-errors cost function, which makes it highly sensitive to outliers [4, §5.4]. To address the sensitivity of MAP-based NLS estimators to outliers, a variety of strategies have been explored, particularly in UWB-based state estimation. For example, Omni-Swarm [17] introduced a visual–inertial–UWB estimation pipeline formulated as a MAP-based NLS optimization problem, where UWB outliers were rejected using fixed elevation-angle criteria. Distributed approaches [7, 8, 6] extend this concept to multi-agent settings, also solving MAP NLS problems while rejecting inconsistent range measurements through averaging, geometric or max velocity constraints. More recently, learning-based methods have been employed to detect and discard NLOS UWB measurements. In [13] and [9], convolutional neural network (CNN) and long short-term memory (LSTM) architectures were proposed to classify and reject suspected NLOS observations. However, since no model can guarantee perfect outlier identification, residual contamination typically persists in the dataset, introducing skewed measurement noise.
To more systematically account for these residual effects, M-estimators incorporate robust cost functions directly within the NLS formulation, allowing the optimization to down-weight the influence of large residuals rather than reject measurements outright. Various robust loss functions have been proposed [19], and in the context of UWB localization, several noise parameterizations have been studied. For example, in [18] a Huber loss was integrated into the NLS optimization, while [15] proposed a Cauchy loss, however these symmetric losses ignore the asymmetry present in NLOS noise. To capture skewed characteristics, [Kok_Hol_Schön_2015, 10] used an asymmetric two-piece distribution. A similar asymmetric Cauchy distribution is used in this letter defined by
@@
where $r(\cdot,\cdot)$ is the residual, $c^{-}$ and $c^{+}$ are scale parameters, and $\alpha=2/[\pi(c^{+}+c^{-})]$ normalizes the distribution. The asymmetric Cauchy loss provides outlier robustness while explicitly modeling skewed measurement noise. A secondary flexible alternative for modeling skewed distributions is to employ a Gaussian mixture measurement model [20],
@@
with weights $w_{j}\geq 0,\ \sum_{j}w_{j}=1$ and $\bm{\psi}_{j}=(\mu_{j},\Sigma_{j})$ [12]. Component weights and parameters are calibrated separately, or solved for in parallel with the state $\mbf{x}$, in an Expectation Maximization (EM) framework [5, §9.3]. The Gaussian mixture measurement model of (10) is incorporated into the joint distribution of (7), and MAP estimation (8) produces a state estimate. Although Gaussian mixture models (GMMs) are flexible enough to capture skewed noise distributions, they require additional parameters and, while incorporation into a NLS framework is possible, it demands careful handling [12]. Moreover, a MAP estimator provides only a point estimate of the state, with uncertainty obtained using an ad-hoc Laplace approximation [4, §4.3.1]. Alternatively, while the methods discussed thus far rely on MAP-based optimization, ESGVI offers a fundamentally different probabilistic framework that estimates both the state and its uncertainty directly through variational inference. Performing parameter learning in parallel with ESGVI [16] has been shown to improve robustness, as jointly estimating the covariance with an appropriate prior implicitly rejects outliers. While robustness was identified as an open direction in [2], measurement models specific to a non-Gaussian distribution, such as the Skew-Laplace in (6) have not yet been studied directly within the ESGVI framework. Consequently, in this work, comparisons are made against robust MAP approaches rather than existing ESGVI methods. In the following, ESGVI is shown to naturally incorporate robust, non-Gaussian measurement models with matrix Lie group states, enabling principled estimation under heavy-tailed or asymmetric noise. This is illustrated using a Skew-Laplace model for UWB localization, capturing NLOS-induced asymmetry.
@@ IV Gaussian Variational Inference
As introduced previously, an alternative to MAP estimation is Gaussian variational inference (GVI), which fits a Gaussian approximation $q(\mbf{x})=\mathcal{N}(\bm{\mu},\bm{\Sigma})$, to the posterior $p(\mbf{x|y,u})$, by minimizing the Kullback-Leibler (KL) divergence [5],
@@
where $\phi(\mbf{x})=-\ln p(\mbf{x,y,u})$ is the negative log joint state and measurement likelihood, and $|\bm{\Sigma}^{-1}|$ is the determinant of the information matrix of the variational approximation, $q(\mbf{x})$. In the work of [2], an exactly sparse Gaussian variational inference (ESGVI) optimization scheme was defined for the vector state $\mbf{x}$, minimizing $V(q)$. The iterative update to the mean $\bm{\mu}$, and information matrix $\bm{\Sigma}^{-1}$, are defined according to
@@
which reduces the computation to smaller, factor-specific subsets. Next, using Stein’s lemma [4, §2.2.16], it was demonstrated how taking derivatives of $\phi_{k}(\cdot)$ can be avoided [2],
@@
Lastly, these marginal expectations can be approximated using a cubature method [14, §6.3], that is
@@ IV-A ESGVI with Matrix Lie Group States
Up to this point, the discussion has been restricted to vector-valued states $\mbf{x}_{k}$, for simplicity. However, as introduced previously, some robotics states are more naturally expressed as MLG states $\mbf{X}_{k}$. As such the variational algorithm (13) must be reformulated to work with MLG states. To do so, uncertainty must be defined in a way that is compatible with the underlying group structure [3]. A Gaussian distribution associated with a MLG state is defined by perturbing the mean element $\bar{\mbf{X}}$ with a zero-mean Gaussian random variable $\delta\bm{\xi}$. Formally,
@@
Although the expressions appear structurally similar to their Euclidean counterparts, their interpretation on MLGs requires further clarification, provided in Appendix A-B. Numerically, the expectations of (18) are evaluated using a cubature method [14, §6.3] such that
@@
The flexibility of $\phi_{k}(\mbf{X}_{k})$ to represent any likelihood model, including Huber or Cauchy, was noted in [2], and such factors are incorporated into ESGVI using the same procedure described above. Bringing these components together, the generalized form of (13) for MLGs is reformulated as,
@@ V Simulation and Experimental Results
The proposed method is validated in simulation and in experiments, and compared against two robust on-manifold MAP baselines, one employing the asymmetric Cauchy loss (MAP-C) shown in (9) and the other a GMM (MAP-GMM) to model skewed measurements. All estimators are implemented in Python within a common factor-graph-based framework. No comparison is made against a vector-valued ESGVI formulation, as the advantages of MLG-based state representations, such as improved consistency and avoidance of singularities, are well established in the literature [11, 3]. The following sections present results first on simulated data, which allows controlled evaluation of performance, and then on real-world experiments to demonstrate practical applicability.
@@ V-A Simulation Results
where $\hat{\mbf{X}}_{i,k}$, and $\mbf{X}_{i,k}$ are the estimated and ground-truth states, for trajectory $i$, at time $t_{k}$, with RMSE reported separately for rotation ($\mathrm{r}\mathrm{a}\mathrm{d}$) and translation ($\mathrm{m}$). Consistency, meaning the ability of the estimator to correctly quantify its own uncertainty, is measured using the average normalized estimation error squared (aNEES) [1, §5.4],
@@
where a perfectly consistent estimator has an aNEES equal to one, due to the normalization in (28) by the dimension of the state $n_{x}$. Performance comparisons are based on $50$ Monte Carlo trials over a trajectory of $400$ poses, with measurements and initial conditions regenerated randomly in each trial, with odometry being produced at $100$ ($\mathrm{H}\mathrm{z}$), and range measurements occurring at $10$ ($\mathrm{H}\mathrm{z}$). In all trials, the same GMM and Skew-Laplace parameters are used, fit using 5,0005{,}000 samples as shown in Figure 1(b). All algorithms are initialized with dead-reckoned gyroscope and wheel odometry measurements, though ESGVI can optionally be “warm-started” using a MAP-based solution to accelerate convergence. Finally, throughout this section, ESGVI uses a third-order Gauss-Hermite cubature method [14, §6.3] to evaluate the expectations in (19). As seen in Table I, all three estimators have similar performance. The GMM model is the most consistent, while the RMSE is comparable between the ESGVI and the MAP estimators. ESGVI is slightly overconfident compared to the MAP estimators, which follows the results of [5] as minimizing the KL-divergence, $\textrm{KL}(q||p)$ can result in a Gaussian that is too confident, underestimating the variance.
@@ V-B Experimental Setup
Experiments were conducted using the Clearpath Husky vehicle, equipped with four DWM1000 Decawave UWB tags, as shown in Figure 2. These tags measure ranges to a set of fixed UWB anchors positioned throughout the environment, which are then used for localization. UWB ranging relies on measuring the time-of-flight of radio signals [15] and is therefore sensitive to non-line-of-sight (NLOS) propagation and multipath effects, which can introduce bias in the measured ranges, as illustrated in Figure 1. Such conditions are common in indoor environments, where obstacles, clutter, and frequent foot traffic create dynamic obstructions that lead to skewed measurements. To reproduce these effects in a controlled manner, obstacles made of metal, wood, and foam were placed near selected UWB anchors, as shown in Figure 2.
@@
The Clearpath Husky was driven along ten independent trajectories designed to randomly expose the UWB range measurements to alternating LOS and NLOS conditions. An example of the sixth trajectory is shown in Figure 3. Ground-truth poses were recorded at $120$ ($\mathrm{H}\mathrm{z}$) using a Vicon motion capture system and used for estimator performance evaluation. The robot was operated at a nominal forward velocity of $0.4$ ($\mathrm{m}\mathrm{/}\mathrm{s}$), resulting in trajectories of approximately $60$ seconds each in duration within a $40$ ($\mathrm{m}^{2}$) indoor test area. UWB range measurements were collected at $30$ ($\mathrm{H}\mathrm{z}$) per tag, yielding over 75,00075{,}000 range observations across all ten trajectories. In addition to NLOS effects, UWB measurements are affected by other error sources, including antenna delays, irregular radiation patterns, and clock skews between UWB tags. To account for these factors, the calibration procedure described in [15] was applied. In particular, the antenna-delay calibration was performed, and the data-driven model relating received signal power to attitude-dependent range bias was reused to isolate for NLOS-induced bias effects. Range measurement errors were then computed with respect to the ground-truth poses, across all trajectories. The resulting error distribution, shown in Figure 4, clearly demonstrates the influence of obstacles through its pronounced skewed tails. Based on this distribution, the asymmetric Cauchy, GMM, and Skew-Laplace models were fitted. The GMM utilized three mixture components, consistent with the methodology of [20], where UWB errors were modeled in a similar fashion. All three models effectively capture the skewness arising from the prevalence of NLOS measurements, with the Skew-Laplace model exhibiting the most conservative behavior for highly biased ranges, while the asymmetric Cauchy and GMM models more closely align with the empirical histogram. Notably, the experimental error distribution closely mirrors the simulated case shown in Figure 1(b), reinforcing the validity of the simulation assumptions.